# takeawaybox backgroundcolor=gray!30, shadow=false, innerleftmargin=5pt, innerrightmargin=5pt, roundcorner=10pt, skipabove=12pt

### v3

@@ 10.2.1 Prompting
Changing the (system-)prompt given to an LLM is the most straightforward and commonly used way to steer the behavior of a model. However, it does so without any changes to weights or activations and does not provide any interpretability benefits. Furthermore, RepE does not take up space in the context window, while the additional tokens from prompting, especially when using many in-context examples, increases the computational cost. Similar to RepE, soft-prompting modifies the model in a continuous embedding space. It does so by operating on the input embeddings, whereas RepE modifies internal activations or weights. However, soft-prompting does not have the goal of identifying the representation of a concept.
Empirical Comparison. There are 24 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a prompting method. Across the surveyed papers, RepE is more effective than prompting at steering the target concept in $75\%$ of the cases (see Table 11). Furthermore, it often had less impact on the general capabilities of the model. Notably, AxBench (Wu et al., 2025) conducts a thorough comparison between multiple RepE methods and prompting and finds prompting to perform better at detecting and steering concrete content-related concepts.
@@
Control Effectiveness Capability Retention Control Effectiveness Capability Retention Control Effectiveness Capability Retention
@@ 10.2.2 Fine-tuning
Fine-tuning modifies the weights of a pre-trained model by further training it on a specific dataset or reward signal. While fine-tuning does influence the model’s representations, it does not do so in a manner that is targeted to identifying and controlling representations of specific concepts. Instead, it applies some uninterpretable, non-sparse changes to the weights that lead to the desired outputs. Unlike RepE, fine-tuning does not allow for precise and isolated control of concepts of interest. Furthermore, many RepE methods provide continuous control, where a concept can be increased or decreased to the desired intensity. As mentioned in Section 6.2.1, there is an overlap between fine-tuning and RepE methods that do output optimization and apply a weight steering function.
Empirical Comparison. There are 25 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a fine-tuning method. Hereby RepE methods such as ReFT (Wu et al., 2024c) and BIPO Cao et al. (2024a) that are adjacent to fine-tuning are counted towards RepE. Table 11 shows that RepE steers the target concept more effectively in $72\%$ of comparisons, while fine-tuning is only better in $20\%$ of cases. A majority of comparisons find that fine-tuning deteriorates general model capabilities more heavily in most cases. Figure 3 lightly indicates that RepE might be more sample effective, outperforming fine-tuning in every comparison under $500$ samples.
@@ 10.2.3 Decoding-based Methods
Decoding-based approaches steer the behavior of LLMs by shifting the token probabilities during the decoding step. Contrastive Decoding takes two sets of token probabilities, contrasts them against each other, and consequently shifts the token probabilities (Li et al., 2023b). Often, this means taking predictions from two different models, one of which is better at the desired behavior. Alternatively, contrastive predictions can be taken from the same model on contrastive inputs.
Similar to RepE methods, this method steers the behavior of the model during inference time. However, contrastive decoding requires multiple forward passes per generated token. Furthermore, it only superficially shifts token probabilities and does not tap into the internal representations of LLMs.
Empirical Comparison. There are 7 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a decoding-based method. RepE-based methods outperform decoding-based methods in $71\%$ of cases and have much less impact on general model capabilities.
@@ 10.2.4 Combining Methods
{mdframed}
[style=takeawaybox] Takeaway: Comparing RepE to Other Methods Other methods for identifying and editing concept representations, like Activation Patching or Concept Erasure, are less applicable for controlling output behaviors or have some overlap with RepE. Our meta-survey shows that RepE offers more effective control at a lower decrease of capabilities compared to prompting, fine-tuning, and decoding-based approaches.
@@ 11 Why Does It Work?
{mdframed}
[style=takeawaybox] Takeaway: Why it works Models represent human-understandable concepts in structured, possibly linear ways. By manipulating these latent variables, we can influence the model’s later computations. LLMs remain functional since they are somewhat resilient to internal manipulations.
@@ 12.1 Empirical Weaknesses
Multi-Concept Control. Applying RepE for multiple concepts at once reduces the effectiveness of steering (van der Weij et al., 2024; Scalena et al., 2024) and leads to larger reductions in model capabilities (Scalena et al., 2024). This might be because applying multiple edits can lead to interference between concept representations (van der Weij et al., 2024). An exception is that combining specific vectors derived by BiPO Cao et al. (2024a) seems to increase their effectiveness. Improving multi-concept control is recognised as one of the major challenges in RepE (Cao et al., 2024a; Postmus & Abreu, 2024; van der Weij et al., 2024; Dong et al., 2024b; Chu et al., 2024). While there are proposals to steer different concepts in different layers (van der Weij et al., 2024), this limits the number of concepts that can be steered simultaneously. Furthermore, Scalena et al. (2024) dynamically adjust the strength of interventions which reduces the negative effects of multi-concept control. Further, there are attempts to control broader representations, like “human preferences”, that encapsulate multiple desired concepts Liu et al. (2023). Nguyen et al. (2025) improve multi-concept steering by sparsely applying each steering vector only at some tokens and enforcing orthogonality between steering vectors to reduce interference. Lastly, Beaglehole et al. (2025) report progress on multi-concept steering based on non-linear feature learning.
@@
Long-Form Generation. While RepE can effectively control short answers, there is a lack of evidence for its ability to control long-form generations and multi-turn conversations. Indeed, there is preliminary evidence that operators identified from a training set with short answers do not effectively steer long generations (Pres et al., 2024). This calls for a thorough evaluation of new RepE methods for longer generations. Recent work has started to address this problem by retaining the natural distribution of activations (Cao et al., 2024a) and dynamically adjusting the representation control throughout a generation (Scalena et al., 2024).
@@
Out-of-Distribution Generalization. To effectively steer a concept, our concept operator should generalize from its training data to different contexts, prompt templates, or different generation tasks. Additionally, if a concept operator does not generalize across different situations, we cannot argue that we have found a general representation of the concept. However, this still presents a challenge for current RepE pipelines, which become less effective when the system prompt differs between the training and test set (Tan et al., 2024) and when the operator is identified after the answer but employed before the answer (Simhi et al., 2024). Evaluating and improving the ability of RepE methods for OOD generalization will be crucial (Mack & Turner, 2024a; Zou et al., 2023a; Li et al., 2023a; Ackerman, 2024).
@@
Deterioration of Capabilities. RepE generally leads to a reduction in general language modeling capabilities and the quality and diversity of generated text (Scalena et al., 2024; Stickland et al., 2024; Zhang et al., 2024c). Aside from the quality of the generated text, Park et al. (2024a) points out that steered models can be worse at instruction following. While this reduction is rather small, it still represents an important cost to employing RepE that should be reduced. Increasing the strength of the intervention reduces the capabilities (Wu et al., 2025; Durmus et al., 2024). This effect could appear because RepE disrupts the underlying structure of a model’s activations (Zhang et al., 2024c). Thus, finding improvements to reduce or better navigate the trade-off between control effectiveness and general model capabilities will be crucial (Li et al., 2024b; Wu et al., 2024b; Wang et al., 2024c). Efforts to solve this problem include fine-tuning the model to prevent reductions in quality through RepE (Stickland et al., 2024), moderating the intervention strength to only steer when necessary (Li et al., 2024d), and more precisely localising the concept representations (Wang et al., 2024b). Others attempt to mitigate this by retaining the underlying structure of the LLM (Zhang et al., 2024c), keeping the magnitude of activations consistent (Pham & Nguyen, 2024), or providing an output-based goal to retain text quality (Xu et al., 2024b).
@@
Learning Specific Concept Representations. Ideally, RepE would steer a concept without influencing other, unrelated concepts. However, in practice, RepE methods struggle to isolate a concept representation from other concepts. For example, steering towards happiness reduces refusal rates for harmful requests (Zou et al., 2023a), steering away from harmfulness makes the model less helpful (Wolf et al., 2024), and steering gender awareness can increase age-based biases (Durmus et al., 2024).
@@
Learning Complete Concept Representations. We would like RepE to identify a concept operator which represents the concepts with all its aspects and in different contexts. A concept like Honesty is multi-faceted and is applied differently depending on the context. However, identifying such complex representations is challenging and likely requires a large coverage of aspects and contexts in the training data, as well as a concept operator that is expressive enough. Chen et al. (2024b) address this by learning multiple orthogonal steering vectors that are supposed to capture different aspects of a concept.
@@
Unreliability. RepE methods are generally unreliable and tend to fail in three ways. Firstly, they are sensitive to the hyperparameters chosen. Often, a small seemingly insignificant change in hyperparameters can cause significant performance decreases (Zhang et al., 2024c; Adila et al., 2024b). This indicates that the chosen RepE setup is not robust. Secondly, there are often concepts that cannot be successfully steered. While it is common for papers to mainly report on the successful applications of their method, more neutral investigations often find concepts that can not be successfully steered (Tan et al., 2024; Tanneru et al., 2024). Braun et al. (2025) provide evidence that some concepts are not effectively steered because the difference between positive and negative activations is not consistent and not linearly separable. Lastly, even for interventions that are successful in steering the average behavior, there are inputs for which the steering effect is negative (Tan et al., 2024; Stickland et al., 2024). This can especially be the case for examples that were already positive with respect to the concept (Im & Li, 2025). This unreliability hinders the deployment of RepE methods.
@@
Unreliable Interpretability. RepE is often used as a tool to interpret LLM representations, but it is questionable whether they provide strong evidence or are largely misleading. Firstly, a failure to identify an effective linear operator for a concept does not prove that the concept is not linearly represented in the model. It could merely be a failure of the method to find the relevant directions. Secondly, it has been shown that finding an effective intervention in specific layers or attention heads does not prove that the concept is actually localized there (Tan et al., 2024). Thirdly, there can be multiple different vectors that successfully steer the concept. Consequently, a direction that steers the model is not necessarily the single direction that represents the concept (Subramani et al., 2022).
@@ 12.2.1 Challenges in Representation Identification
Spuriously Correlated Concepts. Current RI methods are mostly not able to disentangle concepts that are correlated. As a result, the identified concept operator will also capture and steer the spuriously correlated concept (Tan et al., 2024). Concepts can be correlated because they often occur together in the inputs, because the output scoring evaluates both highly or because their representations tend to be activated at the same time. This presents a major hurdle for precisely identifying a concept representation. Sparse Autoencoders might be able to address this by identifying disentangling concept representations.
Concept Misspecification. Representation Identification forces us to specify the target concept by providing a scoring function or inputs related to the concept. Through these, we are hoping to elicit the model’s representation of this concept. However, if the concept specification differs from the intended concepts, the identified concept operator will not be accurate. Deng et al. (2025) point out that RI methods based on contrastive pre-prompts implicitly assume that the model actually follows the pre-prompt.
Furthermore, even when the specified inputs or scoring functions are accurate, they might activate other representations than the ones for the desired concept. For example, we might instruct a model to be honest or dishonest and hope this activates the model representations of “I will be honest/dishonest”. However, for a model that was trying to deceive, such an input might instead activate a representation for “The human wants me to be honest/dishonest”, which is importantly different from the intended concept.
@@
Interference from Superposition. Interpretability research has shown that LLMs are polysemantic since they represent more features than they have neurons (Elhage et al., 2022). Thus, networks represent features in superposition, meaning that features are not all orthogonal to each other, which results in interference between features. In practice, this will mean that controlling a concept representation will also steer some other concepts. Thus, concept operators are not specific to a concept and have side effects on unrelated concepts. Nguyen et al. (2025) attempt to reduce interference by enforcing orthogonality between concept vectors, thus improving multi-concept steering.
Assumptions on Available Data. RepE methods often only work when specific data is available. Firstly, it is not always possible to provide contrastive inputs that describe the behavior and its opposite (Jorgensen et al., 2024; Cai et al., 2024; Postmus & Abreu, 2024). Secondly, many methods rely on expensive human-annotated ground-truth data (Adila et al., 2024b; a). Lastly, methods are not designed to handle noisy and biased data Adila et al. (2024a).
Reliance on Models’ Own Representations. RepE methods tap into the existing representations of a model. Consequently, a concept cannot be controlled with RepE if the network did not learn a representation for it.
Interpretability Lacks Ground Truth. The challenge of developing methods that find representations related to a concept is very difficult because there is no ground truth for where and how the model represents a concept. Thus, we cannot know whether we have identified an accurate representation or what the semantic significance of the concept operator is (Herrmann & Levinstein, 2024).

### v4

@@ 11.2.1 Prompting
There are 24 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a prompting method. Across the surveyed papers, RepE is more effective than prompting at steering the target concept in $75\%$ of the cases (see Table 13). Furthermore, it often had less impact on the general capabilities of the model. Notably, AxBench (Wu et al., 2025) conducts a thorough comparison between multiple RepE methods and prompting and finds prompting to perform better at detecting and steering concrete content-related concepts.
@@
Control
Effectiveness
Capability
Retention
Control
Effectiveness
Capability
Retention
Control
Effectiveness
Capability
Retention
@@ 11.2.2 Fine-tuning
There are 25 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a fine-tuning method. Hereby RepE methods such as ReFT (Wu et al., 2024c) and BIPO Cao et al. (2024a) that are adjacent to fine-tuning are counted towards RepE. Table 13 shows that RepE steers the target concept more effectively in $72\%$ of comparisons, while fine-tuning is only better in $20\%$ of cases. A majority of comparisons find that fine-tuning deteriorates general model capabilities more heavily in most cases. Figure 3 lightly indicates that RepE might be more sample effective, outperforming fine-tuning in every comparison under $500$ samples.
@@ 11.2.3 Decoding-based Methods
There are 7 papers (listed in Appendix D) that provide experiments comparing the effectiveness between RepE and a decoding-based method. RepE-based methods outperform decoding-based methods in $71\%$ of cases and have much less impact on general model capabilities.
@@ 11.2.4 Combining Methods
Takeaway: Comparing RepE to Other Methods
Other methods for identifying and editing concept representations, like Activation Patching or Concept Erasure, are less applicable for controlling output behaviors or have some overlap with RepE. Our meta-survey shows that RepE offers more effective control at a lower decrease of capabilities compared to prompting, fine-tuning, and decoding-based approaches.
@@ 12 Why Does It Work?
Takeaway: Why it works
Models represent human-understandable concepts in structured, possibly linear ways. By manipulating these latent variables, we can influence the model’s later computations. LLMs remain functional since they are somewhat resilient to internal manipulations.
@@ 13.1 Empirical Weaknesses
Multi-Concept Control. Applying RepE for multiple concepts at once reduces the effectiveness of steering (van der Weij et al., 2024; Scalena et al., 2024) and leads to larger reductions in model capabilities (Scalena et al., 2024). This might be because applying multiple edits can lead to interference between concept representations (van der Weij et al., 2024). An exception is that combining specific vectors derived by BiPO Cao et al. (2024a) seems to increase their effectiveness. Improving multi-concept control is recognised as one of the major challenges in RepE (Cao et al., 2024a; Postmus & Abreu, 2024; van der Weij et al., 2024; Dong et al., 2024b; Chu et al., 2024). Multi-property steering is important since model providers will want to steer for more than one concept at a time. For example when van der Weij et al. (2024) combine steering vectors for Myopia, Wealth Seeking, Sycophancy, Agreeableness and Anti-Immigration together, the steering effect for each concept is reduced.
While there are proposals to steer different concepts in different layers (van der Weij et al., 2024), this limits the number of concepts that can be steered simultaneously. Furthermore, Scalena et al. (2024) dynamically adjust the strength of interventions which reduces the negative effects of multi-concept control. Further, there are attempts to control broader representations, like “human preferences”, that encapsulate multiple desired concepts Liu et al. (2023). Nguyen et al. (2025) improve multi-concept steering by sparsely applying each steering vector only at some tokens and enforcing orthogonality between steering vectors to reduce interference. Lastly, Beaglehole et al. (2025) report progress on multi-concept steering based on non-linear feature learning.
@@
Long-Form Generation. While RepE can effectively control short answers, there is a lack of evidence for its ability to control long-form generations and multi-turn conversations. It would be problematic if RepE would lose it’s ability to control LLM behavior when a long text is generated or the conversation continues for multiple turns. Indeed, there is preliminary evidence that operators identified from a training set with short answers do not effectively steer long generations (Pres et al., 2024). This calls for a thorough evaluation of new RepE methods for longer generations. Recent work has started to address this problem by retaining the natural distribution of activations (Cao et al., 2024a) and dynamically adjusting the representation control throughout a generation (Scalena et al., 2024).
@@
Out-of-Distribution Generalization. To effectively steer a concept, our concept operator should generalize from its training data to different contexts, prompt templates, or different generation tasks. For example a concept operator for harmfulness might only be derived on english inputs and fail to prevent harmful responses in other languages. Additionally, if a concept operator does not generalize across different situations, we cannot argue that we have found a general representation of the concept. However, this still presents a challenge for current RepE pipelines, which become less effective when the system prompt differs between the training and test set (Tan et al., 2024) and when the operator is identified after the answer but employed before the answer (Simhi et al., 2024). Evaluating and improving the ability of RepE methods for OOD generalization will be crucial (Mack & Turner, 2024a; Zou et al., 2023a; Li et al., 2023a; Ackerman, 2024).
@@
Deterioration of Capabilities. RepE generally leads to a reduction in general language modeling capabilities and the quality and diversity of generated text (Scalena et al., 2024; Stickland et al., 2024; Zhang et al., 2024c). Aside from the quality of the generated text, Park et al. (2024a) points out that steered models can be worse at instruction following. While this reduction is rather small, it still represents an important cost that might stop model providers from using RepE. Increasing the strength of the intervention reduces the capabilities (Wu et al., 2025; Durmus et al., 2024). This effect could appear because RepE disrupts the underlying structure of a model’s activations (Zhang et al., 2024c). Thus, finding improvements to reduce or better navigate the trade-off between control effectiveness and general model capabilities will be crucial (Li et al., 2024b; Wu et al., 2024b; Wang et al., 2024c). Efforts to solve this problem include fine-tuning the model to prevent reductions in quality through RepE (Stickland et al., 2024), moderating the intervention strength to only steer when necessary (Li et al., 2024d), and more precisely localising the concept representations (Wang et al., 2024b). Others attempt to mitigate this by retaining the underlying structure of the LLM (Zhang et al., 2024c), keeping the magnitude of activations consistent (Pham & Nguyen, 2024), or providing an output-based goal to retain text quality (Xu et al., 2024b).
@@
Learning Specific Concept Representations. Ideally, RepE would steer a concept without influencing other, unrelated concepts. However, in practice, RepE methods struggle to isolate a concept representation from other concepts. For example, steering towards happiness reduces refusal rates for harmful requests (Zou et al., 2023a), steering away from harmfulness makes the model less helpful (Wolf et al., 2024), and steering gender awareness can increase age-based biases (Durmus et al., 2024). A failure to isolate a specific concept representation might cause undesired side-effects that make RepE less usable in practice.
@@
Learning Complete Concept Representations. We would like RepE to identify a concept operator which represents the concepts with all its aspects and in different contexts. A concept like Honesty is multi-faceted and is applied differently depending on the context. Failing to identify a complete representation of honesty means some aspects of the concept are not improved and the steering might fail in some contexts. However, identifying such complex representations is challenging and likely requires a large coverage of aspects and contexts in the training data, as well as a concept operator that is expressive enough. Chen et al. (2024b) address this by learning multiple orthogonal steering vectors that are supposed to capture different aspects of a concept.
@@
Unreliability. RepE methods are generally unreliable and tend to fail in three ways. Firstly, they are sensitive to the hyperparameters chosen. Often, a small seemingly insignificant change in hyperparameters, such as small changes to the steering strength, can cause significant performance decreases (Zhang et al., 2024c; Adila et al., 2024b). This indicates that the chosen RepE setup is not robust. Secondly, there are often concepts that cannot be successfully steered. While it is common for papers to mainly report on the successful applications of their method, more neutral investigations often find concepts, like reducing narcissistic behavior, that can not be successfully steered (Tan et al., 2024; Tanneru et al., 2024). Braun et al. (2025) provide evidence that some concepts are not effectively steered because the difference between positive and negative activations is not consistent and not linearly separable. Lastly, even for interventions that are successful in steering the average behavior, there are inputs for which the steering effect is negative (Tan et al., 2024; Stickland et al., 2024). This can especially be the case for examples that were already positive with respect to the concept (Im & Li, 2025). This unreliability hinders the deployment of RepE methods.
@@
Unreliable Interpretability. RepE is often used as a tool to interpret LLM representations, but it is questionable whether they provide strong evidence or are largely misleading. Firstly, a failure to identify an effective linear operator for a concept does not prove that the concept is not linearly represented in the model. It could merely be a failure of the method to find the relevant directions. Secondly, it has been shown that finding an effective intervention in specific layers or attention heads does not prove that the concept is actually localized there (Wang & Veitch, 2024). Thirdly, there can be multiple different vectors that successfully steer the concept. For example, (Goldman-Wetzler & Turner, 2024) find 800 orthogonal vectors that steer “write code”. Consequently, a direction that steers the model is not necessarily the single direction that represents the concept (Subramani et al., 2022).
@@ 13.2.1 Challenges in Representation Identification
Spuriously Correlated Concepts. Current RI methods are mostly not able to disentangle concepts that are correlated. As a result, the identified concept operator will also capture and steer the spuriously correlated concept (Tan et al., 2024). Concepts can be correlated because they often occur together in the inputs, because the output scoring evaluates both highly or because their representations tend to be activated at the same time. For example, instructing a model to write faulty code can simultaneously trigger the representation to write faulty code and the representation for general unethical behavior (Betley et al., 2025; Soligo et al., 2025). Using RI techniques like difference-in-means or linear probes will consequently lead to a steering vector that captures the representation for faulty code and general unethical behavior. This presents a major hurdle for precisely identifying a concept representation. Sparse Autoencoders might be able to address this by identifying disentangling concept representations.
Concept Misspecification. Representation Identification forces us to specify the target concept by providing a scoring function or inputs related to the concept. Through these, we are hoping to elicit the model’s representation of this concept. However, if the concept specification differs from the intended concepts, the identified concept operator will not be accurate. For example, in Output Optimization we might use an LLM to score how honest outputs are to learn a concept operator. But if the LLM-judge fails to detect some forms of dishonesty, the resulting concept operator will also fail to steer these forms of dishonesty.
Furthermore, even when the specified inputs or scoring functions are accurate, they might activate other representations than the ones for the desired concept. For example, we might instruct a model to be honest or dishonest and hope this activates the model representations of “I will be honest/dishonest”. However, for a model that was trying to deceive, such an input might instead activate a representation for “The human wants me to be honest/dishonest”, which is importantly different from the intended concept. Deng et al. (2025) point out that RI methods based on contrastive pre-prompts implicitly assume that the model actually follows the pre-prompt.
@@
Interference from Superposition. Interpretability research has shown that LLMs are polysemantic since they represent more features than they have dimensions (Elhage et al., 2022). Thus, networks represent features in superposition, meaning that features are not all orthogonal to each other, which results in interference between features. In practice, this will mean that controlling a concept representation will also steer some other concepts. Thus, concept operators are not specific to a concept and have side effects on unrelated concepts. For example steering a model towards academic citations could accidentally also steer it to writing HTTP requests if there is interference between these representations Bricken et al. (2023). Nguyen et al. (2025) attempt to reduce interference by enforcing orthogonality between concept vectors, thus improving multi-concept steering.
Assumptions on Available Data. RepE methods often only work when specific data is available. Firstly, it is not always possible to provide contrastive inputs that describe the behavior and its opposite (Jorgensen et al., 2024; Cai et al., 2024; Postmus & Abreu, 2024). For example using contrastive inputs might not be the right methods for non-binary concepts like the days of the week. Secondly, many methods rely on expensive human-annotated ground-truth data (Adila et al., 2024b; a). For example steering against societal biases might involve humans annotating or writing biased and unbiased texts, which can be prohibitively expensive. Lastly, methods are not designed to handle noisy and biased data Adila et al. (2024a).
Reliance on Models’ Own Representations. RepE methods tap into the existing representations of a model. Consequently, a concept cannot be controlled with RepE if the network did not learn a representation for it. If a model never learned a representation for being honest, RI methods will not be able to find an effective concept operator to make the model more honest.
Interpretability Lacks Ground Truth. The challenge of developing methods that find representations related to a concept is very difficult because there is no ground truth for where and how the model represents a concept. Thus, we cannot know whether we have identified an accurate representation or what the semantic significance of the concept operator is (Herrmann & Levinstein, 2024). While we might find a vector that causes the model to be more honest and detect dishonest generations, we have no guarantee that this maps to how the model actually represents honesty.